Reno Company manufactures part No.498 for use in production line. The


manufacturing costs per unit for 20,000 units of part No.498 are as follows:


Direct materials $6


Direct Manufacturing Labor $30


Variable Manufacturing Overhead $12


Fixed Manufacturing Overhead Allocated $16


Total Manufacturing Costs Per Unit $64


Tray Company has offered to sell 20,000 units of part No.498 to Reno for $60 per unit. Reno


will make the decision to buy the part from Tray if there is an overall savings of at least $25,000


for Reno. If Reno accepts Tray’s offer, $9 per unit of fixed overhead allocated would be totally


eliminated. Furthermore, Reno has determined that the released facilities could be used to save


relevant costs in the manufacture of part No.575. For Reno to achieve an overall saving of


$25,000, how much is the amount of relevant cost that would have to be saved by using the


released facilities in the manufacture of part No.575

1
Expert's answer
2022-01-17T16:23:09-0500

Reno total costs

"64\times20 000=1 280 000"

"60\times20 000=1 200 000"


1 280 000-1200 000=600 000

"9\times20000=180 000"


20 000x-1 200 000=25 000

x=61.25


64-61.25=2.75


2.75 is the amount of relevant cost that would have to be saved by using the

released facilities in the manufacture of part No.575